OpenStudy (anonymous):

find the domain of f(x)^-1 if f(x)= 2/x -10 ?

OpenStudy (anonymous):

f(x) = 2/x - 10 => y = 2/x - 10 => y + 10 = 2/x => x = 2/(y+10) which is defined for all real values of y except y = -10 hence the domain is R - {-10}